In a team of 12 persons, 1/3 are women and 2/3 are men. To obtain a
team with 20% women how many men should be hired?

Found 2 solutions by Jessmh123, melzaren:
Answer by Jessmh123(22)

Since there are 12 people in total, that means there’s 4 women and 8 men. There needs to be 20% women, so because we know that 4 is 20% of 20, that means that there has to be a total of 20 people. If you add 8 men to the already existing 12 workers, you’ll have 20. You’ll end up with 4 women which is 20% of 20, and 16 men which is 80% of 20.
The answer is add 8 men.

Answer by melzaren(3)

A few other ways to solve this:
The quickest way, if you know a value, and the percent of the whole it represents, you can do the following:
# of women: 12 * 1/3 = 4 women
We know that in the end, when the team has all it's members, 4 women will equal 20% of the team.
You can divide 4 by .20 to determine the total number of members on the team. This comes from knowing that having a portion of a whole, and knowing what percent it represents of the whole, you can divide through to get the total number.
4 / .20 = 20
20 total team mates - 4 women, minus the 8 original men from (12 * 2/3 = 8) =
20 - 4 - 8 = 8 men (final answer)
You can verify this by seeing that 20 (the total number of players) * .20 (the percent of the team that is women) = 4.
Another way to do this is:


This comes from knowing that in the end, if the women are 20% of the team, then the men are 80%, so we know that
(number of men) / (entire team) = .80
number of men = 8 + x
total team = 12 + x
(8 + x) / (12 + x) = .80